7. Tempo Electronics, Inc., has an inventory of 5,000 unique electronic chip originally purchased at $2.50 each; their market value is now $ 5 each. The production department has proposed to use these by putting each one together with $6 worth of labor and other materials to produce a wristwatch that would be sold for $10. Should that proposal be implemented? Explain from the viewpoints of economic profit and opportunity costs.
1
Expert's answer
2020-11-17T07:28:36-0500

Option one

sell 5000 chips at "\$5= 5\times5000=\$25,000"

cost to Tempo Electronics ="2.50\times5000=\$12,500"

"profit = 25,000- 12,500 = \$12,500"

option two

sell 5000 units at "\$10 =10\times5000 = \$50,000"

cost of manufacture ("6.0 + 2.50 = 8.50) = 5000\times8.50 = \$42,500"

"profit = 50,000-42,500 = \$7,500"

As profit is greater when sold, so option one is more profitable and should be followed by the firm.